2.Majer Corporation makes a product with the following standard costs: Standard Quantity or Hours 6.7...

1.

Majer Corporation makes a product with the following standard costs: 


Standard Quantity or HoursStandard Price or RateStandard Cost Per Unit
Direct materials6.7 ounces$3.00 per ounce$20.10
Direct labor0.7 hours$18.00 per hour$12.60
Variable overhead0.7 hours$3.00 per hour$2.10
The company reported the following results concerning this product in February

Originally budgeted output5,700 units
Actual output5,300 units
Raw materials used in production30,800 ounces
Actual direct labor-hours7,960 hours
Purchases of raw materials33,200 ounces
Actual price of raw materials$92.90 per ounce
Actual direct labor rate$102.40 per hour
Actual variable overhead rate$4.80 per hour

The company applies variable overhead on the basis of direct labor-hours. The direct materials purchases variance is computed when the materials are purchased. 

The materials quantity variance for February is:


2.

Doogan Corporation makes a product with the following standard costs: 


Standard Quantity or HoursStandard Price or Rate

Direct materials9.2 grams$3.80 per gram

Direct labor0.3 hours$38.00 per hour

Variable overhead0.3 hours$8.80 per hour


The company produced 7,000 units in January using 41,110 grams of direct material and 2,560 direct labor-hours. During the month, the company purchased 46,200 grams of the direct material at $3.50 per gram. The actual direct labor rate was $37.30 per hour and the actual variable overhead rate was $8.60 per hour. 

The company applies variable overhead on the basis of direct labor-hours. The direct materials purchases variance is computed when the materials are purchased. 

The variable overhead rate variance for January is:


3.

Doogan Corporation makes a product with the following standard costs: 







Standard Quantity or  HoursStandard Price or Rate

Direct materials2.0 grams$7.00 per gram

Direct labor0.9 hours$17.00 per hour

Variable overhead0.9 hours$3.00 per hour

The company produced 4,500 units in January using 10,150 grams of direct material and 2,130 direct labor-hours. During the month, the company purchased 10,720 grams of the direct material at $ 7.35 per gram. The actual direct labor rate was $ 17.90 per hour and the actual variable overhead rate was $ 2.80 per hour.


The company applies variable overhead on the basis of direct labor-hours. The direct materials purchases variance is computed when the materials are purchased.

The materials quantity variance for January is:
Answer #1

Post only one question. Post remaining seperately

1. Materials Quantity Variance

= (SQ-AQ) *SP

= (5300*6.7 - 30,800)*3

= (35,510 - 30,800)*3

= 14,130 F